OpenStudy (anonymous):

for every positive integer n, prove that 1+2+...+n= n(n+1)/2

OpenStudy (anonymous):

this is just a guess but it has some to do with n!

OpenStudy (anonymous):

^thing

OpenStudy (anonymous):

so take for instance n=3! so the left will look like this. 3+2+1=3!(3!+1!)/2

OpenStudy (anonymous):

i know that i have to use mathematical induction but I can't seen to get past (n+1)(n+2)/2

OpenStudy (anonymous):

n!=n(n+1)(n+2)....3+2+1is all i have

OpenStudy (anonymous):

thanks but I am really looking for some kind of induction step. That's what we went over in class

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

I found a webpage with this question here: http://www.purplemath.com/modules/inductn.htm

OpenStudy (anonymous):

\[assume 1+2+3+...+k=\frac{ k \left( k+1 \right) }{ 2 } tobe true\] adding both sides k+1 \[1+2+3+...+k+(k+1)=\frac{ k \left( k+1 \right) }{2 }+\left( k+1 \right)\] \[=\left( k+1 \right)\left( \frac{ k }{2 }+1 \right)=\frac{ \left( k+1 \right)\left( k+2 \right) }{ 2 }\] \[=\frac{ \left( k+1 \right)\left( k+1+1 \right) }{2 }\] which shows if it is true for n=k, it is also true for n=k+1 hence by induction it is true for all n.